# An Adaptive Computational Intelligence Approach to Personalised Health Assessment and Immune Age Characterisation from Common Haematological Markers

### v2

@@ 5 Immune age estimation
Here we present a method where a single blood test, as in a complete full blood count of 13 blood analytes, gives insight into the immune age as compared to biological age.
@@ 5.1 Immune Score Over time
To test the hypothesis, we selected two cohorts comprising all the entries with ages ranging between 20 to 30 years, separating them by sex (male or female). These cohorts were used to build two immune spaces by computing the respective mean and standard deviations for each of the thirteen analytes. Afterward, for each sex, we selected seven cohorts of adults aged 20 to 25, 25 to 35, 35 to 45, 45 to 55, 55 to 65, 65 to 75, and 75 to 85 years of age, making for a total of 14 cohorts. Finally, we computed the immune score for each entry. The mean immune age for each of the cohorts is shown in Fig. 9.
Figure 9: (a) The mean immune age as a function of age for NHANES data. The reference ranges were computed over the population aged 20 to 30 for males and females, respectively. Each of the data points represents the mean immune score for the cohort [age5,age+5][age-5,age+5]; for example, 50 represents the immune score for the population aged 45 to 55 years. (b)The distribution of immune age as a function of age. While a correlation trend is visible especially for males, across sex weak because healthy and non-healthy cohorts (both self-reported and ground truth) are confounded.
The results show significantly different behaviour over time for male and female cohorts. For subjects over the age of 30, it is evident that the male population presents an upward trajectory with respect to age for the mean immune score, as well as an increase in the 50th percentile and variance. However, the trend in females is slowed down (compared to males) between ages 40 and 50 and reversed between ages 50 and 60. It remains to be investigated whether some of the analytes involved are confounded with contradictory trends from other physiological differences, such as menopause, which occurs exactly between the ages where the slow-down and reversal in the age score trend for females is detected. It is known and has been reported that pregnancy and menopause produce a higher and more progressive increase in red cell counts, haematocrit levels and increased mean cell volume (MCV), and haemoglobin concentrations [21, 4, 9, 27, 1].
@@ 5.2 Immune age based on the immune score
As explained in Section 5.1, the data compiled comprises a list of seven values representing the mean immune score for each of the seven age cohorts [^1] The corresponding ages are 20 to 25, 25 to 35, 35 to 45, 45 to 55, 55 to 65, 65 to 75 and 75 to 85.. In order to apply the Definition 5.1, we fitted a linear model of degree 3 to the data, obtaining a monotonously increasing function that allows us to interpolate and extrapolate the expected immune score over specific ages. The resulting curve, which we call an immune age curve, is shown in Fig. 10.
Figure 10: A linear approximation of degree 3 to the mean score over seven age cohorts. The linear model smooths the curve and allows for extrapolation.
We will only consider integer ages in years. Therefore we can represent the immune age curve by a list of 70 real values that contains the inferred immune score for each age between 20 and 90. Let us denote this curve by the list of pairs $\text{IAC}=\langle(20,3.0),(21,3.1),\dots,(a,\alpha),\dots,(90,4.75)\rangle$.
@@
$A(\alpha)=\begin{cases}20\quad&\text{if}\,\alpha\leq 3\\ 90\quad&\text{if}\,\alpha\geq 4.75\\ a.(a,\alpha^{*})\in\text{IAC}\quad&\text{elsewhere}\end{cases}$
@@
Figure 11: The immune age is the age as a function of the expected immune score. The slight distortions are an artifact of the discretisation.
As a tool that returns an estimated age with respect to a score, we can analyse its predictive power. Using the NHANES data and focusing on males aged 20 to 90, we can measure the error of using the score to predict a person’s biological age.
As shown in Figure 14, prediction error rates are much lower in healthy NHANES data, compared to unhealthy groups, for both sexes and therefore can be used to quantify the ages of healthy patients to some extent. Prediction error rates for the immune score age were overestimated by approximately 7 years, for the healthy groups, in both sexes. Thus, the age prediction error rate for healthy people is small, with a mean error rate of 7.33 $\pm$ 1.21 for females, and 7.62 $\pm$ 1.17 for healthy males. Furthermore, there is a statistically significant difference between the healthy and non-healthy groups (regardless of sex) in the error rates of age prediction. Pearson’s correlation statistics on the age error rates predicted, by the health status of the red blood cell count, as discussed before, were an R square value of 0.293 and 0.168, for men (P<< 0.0001) and females (P<<0.013). Age prediction error rates were high and not significant when self-reported health measures were used to separate health status groups, but were significant when taking abnormal CBC values only.
Figure 12: Greatest predictors and separators of age for healthy and unhealthy populations. The abrupt colour change is an artifact of typical normal versus abnormal values. Those in transition through the typical normal interval are missed and only recovered on one side or the other once they have already transitioned. These plots show that these dynamics are strongly predictive of age and health.
Table 5: Statistics of each CBC analyte as a predictor or discriminant of health and age
@@ 6 Conclusion
The numerical score was shown to be able to define an ‘immune age’, with some differences for males and females who are known to show greater fluctuations due to e.g. menopause. In the case of estimation of immune age, the biological age of healthy people was closely predicted by common hematological markers typically measured in FBC or CBC, and for unhealthy self-reported or by typical abnormal reference values, the score was not predictive of biological age, comforming with the theoretical expectation of a good predictor of health by divergence between estimated immune and actual biological age.

### v4

@@ 5 Immune age estimation
Here we present a method where a single blood test, as in a complete full blood count of 13 blood analytes, gives insight into the immune age as compared to chronological age. Chronological age simply indicates the age of the individual, as determined by the length of time someone has lived or is living since their birth. This term should not be confused in the context of other literature, such as age estimates predicted by DNA methylation patterns from the Horvath algorithm.
@@ 5.1 Immune Score Over time
To test the hypothesis, we selected two cohorts comprising all the data entries of individuals with ages ranging between 20 to 90 years, from the dataset, separating them by sex (male or female). These cohorts were used to build two immune spaces by computing the respective mean and standard deviations for each of the thirteen analytes. Afterward, for each sex, we selected seven cohorts of adults aged 20 to 25, 25 to 35, 35 to 45, 45 to 55, 55 to 65, 65 to 75, and 75 to 85 years of age, making for a total of 14 cohorts. Finally, we computed the immune score for each entry. The mean immune age for each of the cohorts is shown in Fig. 9.
Figure 9: (a) The mean immune age as a function of age for NHANES data. The reference ranges were computed over the population aged 20 to 85 for males and females, respectively. Each of the data points represents the mean immune score for the cohort [age5,age+5][age-5,age+5]; for example, 50 represents the immune score for the population aged 45 to 55 years. (b)The distribution of immune age as a function of age. While a correlation trend is visible (especially for males), across sex the correlation is weak because healthy and non-healthy cohorts (both self-reported and by ground truth) are confounded.
The results show significantly different behaviour over time for male and female cohorts. For subjects over the age of 30, it is evident that the male population presents an upward trajectory with respect to age for the mean immune score, as well as an increase in the 50th percentile and variance. The immune score decreases slightly for males between 20 and 25 year old. However, the trend in females is slowed down (compared to males) between ages 40 and 50 and reversed between ages 50 and 60. It remains to be investigated whether some of the analytes involved are confounded with contradictory trends from other physiological differences, such as menopause, which occurs exactly between the ages where the slow-down and reversal in the age score trend for females is detected. It is known and has been reported that pregnancy and menopause produce a higher and more progressive increase in red cell counts, haematocrit levels and increased mean cell volume (MCV), and haemoglobin concentrations [21, 4, 9, 27, 1].
@@ 5.2 Immune age based on the immune score
In a manner akin to that presented in Section 5.1, we compiled data into a list of seven values representing the mean immune score for each of the seven age cohorts [^1] The corresponding ages are 20 to 25, 25 to 35, 35 to 45, 45 to 55, 55 to 65, 65 to 75 and 75 to 85.. In order to apply the Definition 5.1, we fitted a linear model of degree 3 to the data, obtaining a monotonously decreasing function that allows us to interpolate and extrapolate the expected immune score over specific ages. The resulting curve, which we call an immune age curve, is shown in Fig. 10.
Figure 10: A linear approximation of degree 3 to the mean score over seven age cohorts from the healthy groups only. The linear model smooths the curve and allows for extrapolation.
We will only consider integer ages in years. Therefore we can represent the immune age curve by a list of 70 real values that contains the inferred immune score for each age between 30 and 85. Let us denote this curve by the list of pairs $\text{IAC}=\langle(20,6.87),(30,6.91),\dots,(a,\alpha),\dots,(80,5.97)\rangle$.
@@
$A(\alpha)=\begin{cases}20\quad&\text{if}\,\alpha\geq 6.91\\ 85\quad&\text{if}\,\alpha\leq 5.624\\ a.(a,\alpha^{*})\in\text{IAC}\quad&\text{elsewhere}\end{cases}$
@@
As a tool that returns an estimated age with respect to a score, we can analyse its predictive power. Using the NHANES data and focusing on males aged 20 to 90, we can measure the error of using the score to predict a person’s age.
As shown in Figure 14, prediction error rates are much lower in healthy NHANES data, compared to unhealthy groups, for both sexes and therefore can be used to estimate the chronological age of healthy patients. Prediction error rates for the immune score age were overestimated by approximately 7 years, for the healthy groups, in both sexes, allowing a correction (mean error rate of 7.33 $\pm$ 1.21 for healthy females, and 7.62 $\pm$ 1.17 for healthy males). Furthermore, there is a statistically significant difference between the healthy and non-healthy groups (regardless of sex) in the error rates of age prediction. Pearson’s correlation statistics on the age error rates predicted, by the health status of the red blood cell count, as discussed before, were an R square value of 0.293 and 0.168, for men (P<< 0.0001) and females (P<<0.013). Age prediction error rates were high and not significant when self-reported health measures were used to separate health status groups, but were significant when taking abnormal CBC values only.
Table 7: Statistics of each CBC analyte as a predictor or discriminant of health and age.
@@
Table 7 shows the statistics of the predictors and separators of age for healthy and unhealthy populations. These plots show that common blood marker dynamics are predictive of age and health.
There are small but potentially meaningful differences between reference ranges. This could affect the classification of out-of-range counts when applying mismatched country-specific ranges. Using uniform, population-appropriate references is important for generalisable models but our main contribution is to learn each individual personal reference values hence making the reference discrepancies per country less relevant.
@@ 6 Conclusions
The numerical score was shown to be able to define an ‘immune age’, with some differences for males and females who are known to show greater fluctuations due to e.g. menopause. In the case of estimation of immune age, the chronological age of healthy people was closely predicted by common hematological markers typically measured in FBC or CBC, and for unhealthy self-reported or by typical abnormal reference values, the score was not predictive of chronological age, conforming with the theoretical expectation of a good predictor of health by the divergence between estimated immune and actual chronological age.
@@
Acknowledgements
Funding
Santiago Hérnandez-Orozco was supported by grant SECTEI/137/2019 awarded by Subsecretaría de Ciencia, Tecnología e Innovación de la Ciudad de México (SECTEI).
@@ Acknowledgements
All the data was made available through Zenodo at https://zenodo.org/records/10426175